OpenStudy (anonymous):

solve for X (x-1/x^2-5x+6)-(x+1/x^2-4)=3/x62-x-6

OpenStudy (anonymous):

x-1/(x-2)(x-3) - x+1/(x-2)(x+2) =3/ (x-3)(x+2)

OpenStudy (dumbcow):

From here you would combine fractions on left side common denominator is (x-2)(x+2)(x-3) [(x-1)(x+2) - (x+1)(x-3)]/(x-2)(x+2)(x-3) = 3/(x-3)(x+2) Now cross multiply 3(x-2)(x+2)(x-3) = (x-3)(x+2)[(x-1)(x+2) - (x+1)(x-3)] divide by (x+2)(x-3) on both sides, they cancel expand and add like terms 3x - 6 = 3x + 1 which is contradiction so there is no solution check my math, but those are the steps